Prove that the real line is a homeomorphic to the interval (0,1) with the subspace topology


1
Expert's answer
2021-07-19T10:57:06-0400

Let us consider the function "f:(0;1)\to \mathbb{R}" defined as "f(x)=\tan (-\frac{\pi}{2}+\pi x)"

This function is strictly increasing and thus is injective.

We know that "\lim_{x\to 0} f(x) = \lim_{t\to -\pi/2} \tan t = -\infty" and "\lim_{x\to 1} f(x) = \lim_{t\to \pi/2} \tan t = +\infty" and therefore it is surjective.

The function "x\to -\pi/2 +\pi x" is continuous on "(0;1)" and the function "\tan t" is continuous on "(-\pi/2 ;\pi/2)", so "f" is continuous on "(0;1)".

Finally, the inverse function defined as "g(y)= (\arctan(y)+\pi/2)/\pi" is continuous on the real line "\mathbb{R}". Therefore, the function "f" realise a homeomorphism between "(0;1)" and "\mathbb{R}".